OpenStudy (2bornot2b):

\[a\le b+\epsilon\] for all epsilon greater than zero Then \[a\le b\] Am I stating it wrong?

OpenStudy (ash2326):

may or may not be correct, depends on the value of epsilon

OpenStudy (turingtest):

is this a Cauchy thing?

OpenStudy (2bornot2b):

Or consider the following, \[a\le b+\epsilon \] for all epsilon greater than zero. Then \[a<b\] Is this statement correct, or the previous one?

OpenStudy (turingtest):

is that supposed to be if \[a\le b+\epsilon,\epsilon>0\]then\[a<b\]??

OpenStudy (2bornot2b):

that is what I am asking..

OpenStudy (turingtest):

um... yeah, looks like it.

OpenStudy (2bornot2b):

So which statement is correct ? Should it be \[a<b\] or \[a\le b\]

OpenStudy (jamesj):

Your first statement of the problem is correct

OpenStudy (jamesj):

a≤b+ϵ for all epsilon > 0 implies a≤b

OpenStudy (turingtest):

if \[\epsilon>0\]then it should be \[a<b\]if we allowed\[a=b\]adding positive epsilon would make\[a>b\]then we would have\[a<b+\epsilon\]which is not what you wanted

OpenStudy (turingtest):

scratch the a>b, typo

OpenStudy (2bornot2b):

Well, I have a kind of feeling, its as follows. Suppose a is lesser than b for all epsilon added to b, then its obvious that a is lesser than b But if a is equal to b for some small epsilon, then a must become less than b when epsilon is removed

OpenStudy (jamesj):

I've told you the answer. Prove it rigorously. And in a way it's obvious that it can't that the conclusion is a < b. For example if a = b. Then \[ a≤b + \epsilon \] for all \( \epslion > 0 \).

OpenStudy (2bornot2b):

Ok thanks. Got you...

OpenStudy (2bornot2b):

James do you agree with this, that no matter what a and b I take, the relation a=b+ epsilon is never going to happen

OpenStudy (jamesj):

It can't, because if it did for some particular \( \epslion' \), then \[ a > b + \epsilon'/2 \] which violates the hypothesis.

OpenStudy (2bornot2b):

So do you think that the theorem would hold even if it were stated as follows a<b+ϵ for all epsilon > 0 implies a≤b Am I right?

OpenStudy (2bornot2b):

do you think that the theorem would hold even if it were stated as follows a<b+ϵ for all epsilon > 0 implies a≤b Am I right?

OpenStudy (jamesj):

That's correct.

OpenStudy (2bornot2b):

Ok then what is the point of that equality?

OpenStudy (2bornot2b):

The thing that is given is \[a\le b+\epsilon\]

OpenStudy (2bornot2b):

What does that mean? Either \[a< b+\epsilon\] or \[a= b+\epsilon\] implies that \[a\le b\] Am I right?

OpenStudy (jamesj):

There are four possible theorems here 1. \[ (\forall \epsilon > 0) \ a \leq b+\epsilon \implies a \leq b \] 2. \[ (\forall \epsilon > 0) \ a \leq b+\epsilon \implies a < b \] 3. \[ (\forall \epsilon > 0) \ a < b+\epsilon \implies a \leq b \] 4. \[ (\forall \epsilon > 0) \ a < b+\epsilon \implies a < b \] Which of these is true?

OpenStudy (2bornot2b):

I think the first and the third, am I right?

OpenStudy (jamesj):

We're seen why 2 is false. Why is 4 false?

OpenStudy (2bornot2b):

Four is not false, but four is just a particular case of 3, right?

OpenStudy (2bornot2b):

Well my question is the same still Does the theorem means (implicitly ) Either a<b+ϵ or a=b+ϵ implies that a≤b Am I right?

OpenStudy (jamesj):

\[ (\forall \epsilon > 0) \ a \leq b+\epsilon \implies ( (\forall \epsilon > 0) \ a < b+\epsilon) \ or \ ((\forall \epsilon > 0) \ a = b+\epsilon) \] and \[ (\forall \epsilon > 0) \ a \leq b+\epsilon \implies a \leq b \] But that doesn't necessarily mean \[ ( (\forall \epsilon > 0) \ a < b+\epsilon) \ or \ ((\forall \epsilon > 0) \ a = b+\epsilon) \implies a \leq b \] In other words, (A => B) and (A => C) doesn't necessarily mean (B => C). But in this case \[ ( (\forall \epsilon > 0) \ a < b+\epsilon) \ or \ ((\forall \epsilon > 0) \ a = b+\epsilon) \implies ( (\forall \epsilon > 0) \ a < b+\epsilon) \] because the second statement is false. And indeed \[ ( (\forall \epsilon > 0) \ a < b+\epsilon) \implies a < b \implies a \leq b \]

OpenStudy (2bornot2b):

I will not be wasting your time keeping you waiting on this. At present my brain is trying to understand your last post. I will let you know once the processing is over.

OpenStudy (2bornot2b):

But before you leave, I am once again trying to present my question by saying, what is the need of "less than equal to", when the theorem is OK for just a<b+ϵ for all epsilon > 0 implies a≤b

OpenStudy (jamesj):

Actually there's a mistake in my last line \[ (\forall \epsilon>0) a < b+ \epsilon) \implies a \leq b \] but it doesn't imply a < b. Therefore in fact theorem 4 above is false. For example if a = b, then \[ a < b + \epsilon \ \forall \epsilon > 0 \]

OpenStudy (2bornot2b):

A while ago you agreed that no matter what value of a and b I take, I will never get a=b+\epsilon Then what is the point of keeping that equality in the theorem

OpenStudy (jamesj):

What we've shown is \[ (\forall \epsilon>0) a \leq b + \epsilon \implies (\forall \epsilon>0) a < b + \epsilon \] So the equality is irrelevant.

OpenStudy (2bornot2b):

Yes, "equality is irrelevant.", then why keep it?

OpenStudy (jamesj):

because sometimes when you come to use these results in other more important theorems, you'll have an inequality with \[ \leq \] versus with \[ < . \] It's a useful technical thing to know that it doesn't matter; that in both cases you can draw the same conclusion. All of these results are just technical lemmas that are used for more important results. They're also a good training ground for gaining familiarity for these sorts of issues.

OpenStudy (2bornot2b):

Ok then, I will need this later, and now I am satisfied. Now if you don't mind, may I ask, did you find anyone else, before, bothering himself with such a question for such a long time. I mean do you think I was a slow thinker

OpenStudy (jamesj):

The step up in thinking required in introductory analysis to begin manipulating with epsilon-type proofs (like this one; like epsilon-N for sequences and series; like epsilon-delta for continuous functions) is a big step for just about everyone I know. I wouldn't feel bad about it. But do work lots and lots of problems with it so you gain fluency. Just as in HS you didn't "get" quadratic equations the first day you were shown them, but after being forced to solve 100+ of them, you can now do them easily. This is intellectually analogous.

OpenStudy (2bornot2b):

James you don't have the slightest idea how helpful you are. Thanks once again.

OpenStudy (jamesj):

Um thanks. All cash donations gratefully received. ;-)

OpenStudy (2bornot2b):

How to send?

OpenStudy (jamesj):

joke!

OpenStudy (jamesj):

til later

OpenStudy (2bornot2b):

lol

OpenStudy (2bornot2b):

Don't consider this a joke, I am serious, I need your comment on my following statement. \[( (\forall \epsilon > 0) \ a < b+\epsilon) \ or \ ((\forall \epsilon > 0) \ a = b+\epsilon)\ or\ ((\forall \epsilon > 0) \ a = banana)\ or \ ((\forall \epsilon > 0) \ a = grapes) \] \[\implies ( (\forall \epsilon > 0) \ a < b+\epsilon)\implies a\le b\]

OpenStudy (jamesj):

Your equation is chopped off. But the basic principle is A or B or C or D or E => A if B, C, D and E are false.

OpenStudy (2bornot2b):

Here it is (∀ϵ>0) a<b+ϵ) or ((∀ϵ>0) a=b+ϵ) or ((∀ϵ>0) a=banana) or ((∀ϵ>0) a=grapes) =>((∀ϵ>0) a<b+ϵ)=>a≤b

OpenStudy (jamesj):

yes

OpenStudy (jamesj):

because a=banana and a=grapes is false for any value of epsilon

OpenStudy (2bornot2b):

Finally done! Now I am going to write it down somewhere.. Thanks again

OpenStudy (jamesj):

actually, there is one quibble here...

OpenStudy (jamesj):

\[ (\forall \epsilon >0) \ a \leq b + \epsilon \] means that for each such epsilon \[ a < b + \epsilon \] or \[ a = b + \epsilon \] The more logically coherent way to show then that equality is never attained is to imitate the argument we wrote above: If there any particular \( \epsilon > 0 \), call it \( \epsilon' \) such that \[ a = b + \epsilon' \] then it must be the case that \[ a > b + \epsilon'/2 \] which contradicts the hypothesis. Hence \[ (\forall \epsilon >0) \ a \leq b + \epsilon \implies (\forall \epsilon >0) \ a < b + \epsilon \]

OpenStudy (2bornot2b):

Yes, I understand that, but where is the quibble. Is it that, I can't actually show that apple and grape is not possible?

OpenStudy (2bornot2b):

....just like you showed equality is not possible

OpenStudy (jamesj):

No, actually, sigh, I'm embarrassed and annoyed that what I wrote down above isn't correct. One sec ... I'll write it out.

OpenStudy (jamesj):

Consider the following: \[ (\forall n \in \mathbb{N}) ( (n \hbox{ is even}) \ or \ (n \hbox{ is odd})) \] That doesn't imply \[ ((\forall n \in \mathbb{N}) (n \hbox{ is even})) \ or \ ((\forall n \in \mathbb{N}) (n \hbox{ is odd})) \] so far so good?

OpenStudy (2bornot2b):

OK...

OpenStudy (jamesj):

Likewise \[ (\forall \epsilon >0)\ a \leq b + \epsilon \] doesn't imply \[ ( (\forall \epsilon >0)\ a < b + \epsilon ) \ or \ ( (\forall \epsilon >0)\ a = b + \epsilon ) \] If we want to show (and we do!) that \[ (\forall \epsilon >0)\ a \leq b + \epsilon \implies (\forall \epsilon >0)\ a < b + \epsilon \] use the new argument I've just written above. Not the old one, which is wrong.

OpenStudy (jamesj):

The general logical principle here is this: \[ (x \in S) (P(x) \ or \ Q(x)) \] doesn't imply \[ ((x \in S) P(x) ) \ or \ ((x \in S) Q(x)) \]

OpenStudy (2bornot2b):

Yes, I am getting you. Could you please copy paste the part that you wrote wrong previously

OpenStudy (jamesj):

This step is wrong:

OpenStudy (2bornot2b):

OK..

OpenStudy (jamesj):

sorry bout that. (and see? even old(er) dogs make errors.)

OpenStudy (2bornot2b):

LOL

OpenStudy (2bornot2b):

So finally may I say (∀ϵ>0) (a<b+ϵ or a=b+ϵ or a=banana or a=grapes) =>((∀ϵ>0) a<b+ϵ)=>a≤b

OpenStudy (jamesj):

Yes. But reasons in each case are different It can't be the case that a=b+ϵ for any epsilon by the argument above It's not the case that a=banana or a=grapes because a is a number, not a fruit. The general logical principle here then is \[ [(\forall x \in S) (P(x) \lor Q(x))] \land [(\forall x \in S) \lnot Q(x)] \implies (x \in S) P(x) \] where \( \lor \) is or and \( \lnot \) is not